OpenStudy (anonymous):

split 20 into 2 nonnegative numbers x,y such that the product of x and y^2 is a maximum

OpenStudy (anonymous):

\[ x+y = 20 \\ \max(xy^2) \]

OpenStudy (anonymous):

\[ x,y\ge0 \]

OpenStudy (anonymous):

Yes, I found out that the derivative of the primary is (sqrt 5)/(sqrt x) but when solving for x it is not possible?

terenzreignz (terenzreignz):

Hmm... Look at it this way \[\large y = 20 -x\] And find the maximum of \[\large xy^2 = x(20-x)^2=x(x^2-40x+400)\]

OpenStudy (anonymous):

Don't you have to solve for y for the constraint though?

terenzreignz (terenzreignz):

we did, didn't we? And we got y = 20-x

OpenStudy (anonymous):

I believe the constraint is xy^2=20

terenzreignz (terenzreignz):

No, that is what we are supposed to maximise. The constraint is x + y = 20

OpenStudy (anonymous):

Wow I feel stupid, thank you so much!!